OpenStudy (anonymous):

The perimeter of a rectangle field is 380, length is 30 yd longer than the with find the dimensions?

OpenStudy (anonymous):

help anyone?

OpenStudy (anonymous):

Ok. So I have your answer. Now I'll show you how to do it. \[P = 2L + 2W\] Perimeter is equal to 2 sides of length, and 2 sides of width added. So, \[380 = 2(W+30) + 2W\] This is because length is width plus 30, and you add this to the two sides of width. Then, \[380 = 2W + 60 + 2W\] After you multiply everything out. Subtract 60 to the other side, and add the 2W + 2W \[320 = 4W\] Divide by 4, \[W = 80\] And remember that length was W + 30, so \[L = 110\] \[380 = 110+110+80+80\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Of course. Good luck with your hw =).